OpenStudy (anonymous):

A wooden pyramid, 3 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 6 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids.

OpenStudy (campbell_st):

the ratio of lengths is 1:6 the ratio of base area is 1:36 this can be shown if the base side length is 1 unit then Area = 1 unit^2 increasign by a factor of 6 side length 6 Area = 36 units^2 with volume the lnear measurements are cubed 1:6 for volume is 1:216 the volume will be 216 times larger. using a square pyramid base length 1 unit height 3 units V = 1/3 x 1^2 x 3 V = 1 unit^3 increasing all measurements by a factor of 6 V = 1/3x 6^2 x 18 V = 216 units^3